OpenStudy (anonymous):

Find x if f(x) = 2x + 7 and f(x) = -1. -4 -3 5 how do i do this one same concept but its diffrent rite.....@ilfy214

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

ilfy214 you were here first

OpenStudy (anonymous):

If you want to solve it go ahead

OpenStudy (anonymous):

Definitely different because \[f(x) = y = -1\] Now you can plug it in.

OpenStudy (anonymous):

@ilfy214 has been helping me for a while

OpenStudy (anonymous):

Go ahead @Microrobot if you want to! :) We're all helping here

OpenStudy (anonymous):

i dont mind either

OpenStudy (anonymous):

Now you can insert the -1 \[ f(x) = y = -1\]\[ f(x) = 2x + 7\]\[-1 = 2x + 7\]Got it?

OpenStudy (anonymous):

Ok, you can take this one dont worry

OpenStudy (anonymous):

ok so.....lets see here im not getting the way the equation is set can you explainit more????

OpenStudy (anonymous):

i think its 3

OpenStudy (anonymous):

f(x) is also another way of saying y. If -1 is equal to f(x) then it is also equal to y. Look at it like this. Maybe it will help. \[f(x) = -1\]\[y = -1\]\[y = 2x + 7\]\[-1 = 2x + 7\]Can you do the rest?

OpenStudy (anonymous):

@JupitersBrain You gotta explain it first before you just give out the answers.. lolz

OpenStudy (anonymous):

Also, that isn't the answer...

OpenStudy (anonymous):

\[−1=2x+7\]\[−8=2x\]\[−4=x\]Not 3